# Empirical Best Prediction of Poverty Indicators via Nested Error Regression with High Dimensional Parameters

### v1

@@ 5 Uncertainty measures
where $E_{\bm{\phi}}$ denotes expectation with respect to the super-population model. However, the analytical approximations to the MSPE are difficult to derive when involving complex parameters such as the FGT poverty measures.
In this paper, we propose a parametric bootstrap method that can be applied to produce reasonable estimators of various uncertainty measures, not necessarily MSPE. Our bootstrap scheme builds on the method introduced by Lahiri and Salvati (2023), which demonstrated that for known $\tau_{i}$, the parametric bootstrap estimators of uncertainty measures converge in probability to the true corresponding uncertainty measures as $m$ tends to infinity, under certain regularity conditions. Our approach adapts this framework to the context of poverty mapping. The steps of the proposed bootstrap are as follows:
Step 1.
@@
$Y^{*(b)}_{ij}=\hat{\beta}_{0}+\mathbf{x}_{ij}^{\prime}\hat{\bm{\beta}}_{i}+u_{i}^{*(b)}+e^{*(b)}_{ij},$ (20)
where $u_{i}^{*(b)}\sim N(0,\hat{\sigma}_{\gamma}^{2})$ and $e^{*(b)}_{ij}\sim N(0,\hat{\sigma}^{2}_{\epsilon i})$ are all independently distributed, $i=1,\cdots,m$, $j=1,\cdots,N_{i}$.
Step 3.
@@
$Y^{*(b)}_{ij;s}=\hat{\beta}_{0}+\mathbf{x}_{ij;s}^{\prime}\hat{\bm{\beta}}_{i}+u_{i;s}^{*(b)}+e^{*(b)}_{ij;s},i\in s,j=1,\cdots,n_{i},$ (21)
where $\{u_{i;s}^{*(b)};i\in s\}$ are selected using the same area indices as in the original sample from $\{u_{i}^{*(b)};i=1,\cdots,m\}$, which are obtained in Step 2. Here $\{e^{*(b)}_{ij;s};i\in s,j=1,\cdots,n_{i}\}$ are independently generated from $N(0,\hat{\sigma}^{2}_{\epsilon i})$. These errors are not subsets of the population errors generated in Step 2 but are instead independently drawn from the same distributions.
Step 5.
@@
Step 6.
A Monto Carlo approximation to the theoretical bootstrap estimator $\text{MSPE}_{*}(\hat{F}^{EB*(b)}_{\alpha i})=E_{\bm{\phi}^{*}}(\hat{F}^{EB*(b)}_{\alpha i}-F^{*(b)}_{\alpha i})^{2}$ of $\hat{F}^{EB}_{\alpha i}$ is calculated as
@@
6 Design-based Monte Carlo simulations
In this section, we perform design-based simulation experiments to assess the performance of various estimators for poverty measures across repeated samples drawn from a fixed population. Specifically, we utilize the 2021 Italian dataset from the European Survey on Income and Living Conditions (EUSILC), comprising $29,798$ observations. The dataset is treated as the fixed population, with Italy’s $107$ provinces serving as the areas of interest. This setup allows us to compute the true area-level headcount ratios and poverty gaps, serving as a benchmark for evaluation:
$F_{\alpha i}=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}F_{\alpha ij},\quad i=1,\cdots,107,$
where
$F_{\alpha ij}=\bigg{(}\frac{z-E_{ij}}{z}\bigg{)}^{\alpha}I(E_{ij}<z),\;j=1,\cdots,N_{i}.$
Here, $E_{ij}$ denotes the equalized annual net income for each individual, and $z$ represents the fixed poverty line, which is roughly equal to 0.6 times the median of the welfare variables $E_{dj}$ for the above fixed population.
Following Molina and Rao (2010), the welfare variable $E_{ij}$ is transformed to ensure positivity by adding a fixed constant $c$, and then taking the logarithm. Specifically, the response variable is defined as $Y_{ij}=\log(E_{ij}+c)$. The auxiliary variables used in the model comprise both continuous and categorical types. The continuous variables include age, house size in square meters (MQ), household size (hsize), and total years of work experience (Years_work). Binary indicators consist of working status (Work_Status) and gender, while categorical variables include six levels of income and living conditions (cond_ilo) and marital status (Marital_status). Additionally, life satisfaction (life_satisf) is treated as a scale variable.
To explore the data, we calculate the sample variance of the response variable to understand its dispersion within each province. Additionally, we fit a regression model for the response variable using the auxiliary variables as predictors, separately for each province. This approach allows us to examine the relationships between the response and auxiliary variables while accounting for provincial differences in patterns and variability. Figure 2 shows boxplots of the distribution of residuals by province and the distribution of regression coefficients. These results indicate that the regression coefficients vary significantly across provinces, and within-area variation may also differ between provinces. Consequently, the assumptions of the traditional NER model may not hold for this dataset.
(a) Distributions of estimated regression coefficients.
(b) Distributions of the residuals by province.
Figure 2: Distributions of estimated regression coefficients (a) and residuals (b) obtained by fitting a regression model separately for each province.
In the simulation experiments, we consider the following estimators of the FGT measures for each province:
@@ 6 Design-based Monte Carlo simulations
The EBP based on the traditional NER model from Molina and Rao (2010) (MR),
The proposed EBP based on the NERHDP model (CLS),
The ELL estimator using the algorithm from Elbers, Lanjouw and Lanjouw (2003) (ELL).
For the EBP methods, we use $K=100$ for the Monte Carlo approximation, and for the ELL method, 100 simulated censuses are generated.
Two experimental settings are considered: (1) all 107 provinces in the fixed population are sampled; (2) only 90 provinces are sampled, leaving 17 provinces with no observations. For each setting, 1000 samples are drawn using simple random sampling (SRS) without replacement within each sampled area, with the sample size set at $0.1\times N_{i}$. The performance of the above estimators is assessed using the following three criteria:
@@
Relative Bias (RB): The average RB over all small areas. For a given area, RB is defined as the ratio of the average difference between the estimated and true small-area FGT measures to the true small-area FGT measure, averaged over simulations.
@@
Relative Root Mean Squared Prediction Error (RRMSPE): The average RRMSPE over all small areas. For a given area, RRMSPE is the ratio of the square root of the average squared difference between the estimated and true small-area FGT measures to the true measure, averaged over simulations.
Efficiency (EFF) measured as the ratio of the RRMSPE of each estimator/predictor to the RRMSPE of the corresponding the baseline estimator. DIRECT and MR serve as the baseline estimators in experiments 1 and 2, respectively.
First, the computational efficiency of the new parameter estimation algorithm is evaluated, comparing it to the existing algorithm proposed by Lahiri and Salvati (2023). All computations are carried out on a 13th Gen Intel(R) Core(TM) i7-13700F 2.10 GHz processor. For this comparison, five samples are generated using SRS without replacement within each province, with the sample size set at $0.02\times N_{i}$, resulting in a fixed total sample size of 616. Both algorithms are applied to these five samples, and their computation times are recorded. The algorithm proposed by Lahiri and Salvati (2023) requires an average computation time of $1197.59$ seconds, while the new algorithm completes the same task in just $2.17$ seconds on average. This demonstrates a substantial improvement, as our algorithm requires only about $0.2\%$ of the computation time compared to the existing algorithm.
6.1 Experiment 1: All provinces are sampled
In this simulation experiment, we assume that all 107 provinces are included in the samples ($m_{s}=m=107$). For each province, samples are repeatedly drawn with a size equal to $0.1\times N_{i}$. The total sample size is 2979 and the province-specific sample sizes range from 2 to 180. In this scenario, all four estimators are available for all provinces.
Table 2 provides a summary of the overall performance of the four estimators, evaluated based on average RB, average RRMSPE, and EFF, with the DIRECT estimator serving as the baseline for comparison. CLS consistently outperforms MR in terms of both bias and RRMSPE, likely due to the limitations of the MR method, which assumes identical regression coefficients and variance components across the population—assumptions that may not be valid in this context. Additionally, ELL shows better performance than MR, in contrast to the findings in Molina and Rao (2010). This difference may stem from our use of the algorithm from Elbers, Lanjouw and Lanjouw (2003), which incorporates heteroscedasticity at the individual level, rather than the simplified version outlined in Molina and Rao (2010). For the poverty gap indicator, CLS outperforms in both relative bias and RRMSPE, while its performance is comparable to ELL for the headcount ratio.
Table 2: Performance of estimators/predictors of small area FGT measures when all areas are sampled; Average is over 107 areas; numbers in parenthesis are the values of the efficiency over DIRECT in terms of RRMSPE.
Predictor RRMSPE % (EFF) Relative Bias %
Headcount Ratio Poverty Gap Headcount Ratio Poverty Gap
DIRECT 48.71 (1.00) 60.43 (1.00) 4.16 5.14
MR 28.73 (0.59) 33.39 (0.55) 25.32 21.71
CLS 22.94 (0.47) 17.81 (0.29) 15.34 -2.03
ELL 22.48 (0.46) 20.25 (0.34) 15.61 8.18
Figure 3 illustrates the province-specific performance of the estimators. For the poverty gap indicator, Figure 3(a) shows that CLS outperforms MR in terms of both RRMSPE and relative bias for most provinces, with MR exhibiting a tendency toward positive bias. Similarly, Figure 3(b) highlights the performance of the four methods for estimating the headcount ratio, demonstrating that CLS achieves smaller RRMSPE and absolute bias than MR in the majority of provinces.
(a) The RRMSPE and Relative Bias of the four estimators of poverty gap for each area.
(b) The RRMSPE and Relative Bias of the four estimators of headcount ratio for each area.
Figure 3: The RRMSPE and Relative Bias of the four estimators of FGT measures, when all provinces are sampled.
6.2 Experiment 2: 17 out-of-sample provinces
In the second experiment, we assess the performance of EBP methods and the ELL method in estimating FGT measures for out-of-sample provinces, where the direct estimator is unavailable due to the absence of observations. Specifically, 90 provinces are randomly selected, and this selection is fixed to draw simple random samples (SRS) repeatedly from these provinces. Consequently, 17 provinces are consistently excluded from the sampling process, ensuring they have no observations across all simulations.
Table 3 summarizes the overall performance of the three estimators for the out-of-sample provinces. The results align with those observed in Experiment 1. Overall, CLS outperforms MR in terms of both RRMSPE and relative bias for both FGT measures. CLS also shows better performance than ELL for the poverty gap measure, while their performance is comparable for the headcount ratio.

### v2

@@ 5 Uncertainty measures
where $E_{\bm{\phi}}$ denotes expectation with respect to the super-population model (10). However, the analytical approximations to the MSPE are difficult to derive when involving complex parameters such as the FGT poverty measures.
In this paper, we use a parametric bootstrap method that can be applied to produce reasonable estimators of various uncertainty measures, not necessarily MSPE. Our bootstrap scheme builds on the method introduced by Lahiri and Salvati (2023), which is in line with the bootstrap framework introduced by González-Manteiga et al. (2008) and later applied in Molina and Rao (2010), while allowing for a more flexible modeling structure.
The steps of the proposed bootstrap are as follows:
Step 1:
@@
$Y^{*(b)}_{ij}=\hat{\beta}_{0}+\mathbf{x}_{ij}^{\prime}\hat{\bm{\beta}}_{i}+\gamma_{i}^{*(b)}+\epsilon^{*(b)}_{ij}$, (21)
where $\gamma_{i}^{*(b)}\sim\mathcal{N}(0,\hat{\sigma}_{\gamma}^{2})$ and $\epsilon^{*(b)}_{ij}\sim\mathcal{N}(0,\hat{\sigma}^{2}_{\epsilon i})$ are all independently distributed, $i=1,\cdots,m$, $j=1,\cdots,N_{i}$.
Step 3:
@@
$Y^{*(b)}_{ij}=\hat{\beta}_{0}+\mathbf{x}_{ij}^{\prime}\hat{\bm{\beta}}_{i}+\gamma_{i}^{*(b)}+\epsilon^{*(b)}_{ij},i=1,\cdots,m_{s},j=1,\cdots,n_{i},$ (22)
where $\{\gamma_{i}^{*(b)};i=1,\cdots,m_{s}\}$ are selected using the same area indices as in the original sample from $\{\gamma_{i}^{*(b)};i=1,\cdots,m\}$, which are obtained in Step 2. Here $\{\epsilon^{*(b)}_{ij};i=1,\cdots,m_{s},j=1,\cdots,n_{i}\}$ are independently generated from $\mathcal{N}(0,\hat{\sigma}^{2}_{\epsilon i})$. These errors are not subsets of the population errors generated in Step 2 but are instead independently drawn from the same distributions.
Step 5:
@@
Step 6:
A Monto Carlo approximation to the theoretical bootstrap estimator $\text{MSPE}_{*}(\hat{F}^{EB*}_{\alpha i})=E_{\bm{\phi}^{*}}(\hat{F}^{EB*}_{\alpha i}-F^{*}_{\alpha i})^{2}$ of $\hat{F}^{EB}_{\alpha i}$ is calculated as
@@
6 Model-based Monte Carlo simulations
We carry out two model-based simulation studies to evaluate the performance of the proposed EBPs of small area HCRs and PGs. In this paper, we followed Bianchi and Salvati (2015) in choosing $q_{i}$. In the model-based simulations, we use the following two auxiliary variables: $x_{1}$, household size (continuous) and $x_{2}$, land ownership (binary) from the 2001 Albanian census to repeatedly generate synthetic populations. For the synthetic populations, our small areas are communes and we have $m=374$ communes with different subpopulation sizes, $N_{i}$, ranging from 141 to 89764. Then, a sample is drawn from each generated population using a probability sampling design. Additionally, we use different models to generate such synthetic populations for evaluating different predictors under different simulation conditions.
In both simulation studies, we use a simple random sampling from each small area population. Since we focus on predicting small area HCR and PG, the explicit-form of EBP are available. Therefore, we include both the explicit form and Monte Carlo approximation of the EBP under our proposed NERHDP model, which enables us to check the efficiency and variability of Monte Carlo approximation. In addition, we compare our proposed performance to other existing predictors. Specifically, the following predictors of HCR and PG are considered in our model-based simulations:
@@ 6 Model-based Monte Carlo simulations
The Monte Carlo approximation of the EBP based on the traditional NER model from Molina and Rao (2010) (MR),
The explicit-form EBP based on the traditional NER model from Morales et al. (2021) (MRE),
The Monte Carlo approximation of EBP based on the proposed NERHDP model (CLS),
The explicit-form EBP based on the proposed NERHDP model (CLSE),
The simplified ELL-type estimator based on the algorithmic structure of Elbers et al. (2003), implemented under the comparative simulation framework adopted in Molina and Rao (2010) (hereafter, SELL).
Note that both MR and MRE are based on the following NER model:
$y_{ij}=\beta_{0}+\mathbf{x}_{ij}^{\prime}\bm{\beta}+\gamma_{i}+\epsilon_{ij},\quad i=1,\dots,m;\;j=1,\dots,N_{i},$
where $\bm{\beta}$ is the vector of constant regression coefficients, the area effects $\gamma_{i}\sim\mathcal{N}(0,\sigma_{\gamma}^{2})$ defined at the domain level, and the errors $\epsilon_{ij}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ are independent.
The MR predictor for a poverty indicator $F_{\alpha i}$ is approximated using Monte Carlo simulation:
$\hat{F}^{EB}_{\alpha i}=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}\frac{1}{K}\sum_{k=1}^{K}h_{\alpha}\left(y_{ij}^{*(k)}\right),$ (24)
where $y_{ij}^{*(k)}$ is the $k$th simulated value of $y_{ij}$ drawn from its conditional distribution given the sample. In our simulations, we set $K=100$. The MRE approach, using the explicit analytical expressions for the EBP derived under the NER model (6), is described in Chapter 10 of Morales et al. (2021).
While our simulation setup is similar to Molina and Rao (2010) in that clusters are treated as small areas (domains), the simplified ELL-type predictor (SELL) used in our simulation studies adopts the basic algorithmic structure of Elbers et al. (2003), but is implemented under the controlled comparative framework of Molina and Rao (2010). It differs from the usual ELL setup, where clusters are nested within small areas.
Specifically, the underlying model is
$Y_{ij}=\beta_{0}+\mathbf{x}_{ij}^{\prime}\bm{\beta}+\gamma_{i}+\epsilon_{ij},\quad i=1,\dots,m;\;j=1,\dots,N_{i},$ (25)
where the unit-level error term $\epsilon_{ij}$ has a flexible variance structure $\sigma^{2}_{\epsilon ij}$ to accommodate heteroskedasticity. Initial estimates of the regression coefficients in (25) are obtained via OLS estimation. Denote the residuals from this regression by $\hat{r}_{ij}$, which are decomposed as
$\hat{r}_{ij}=\hat{r}_{i.}+(\hat{r}_{ij}-\hat{r}_{i.})=\hat{\gamma}_{i}+\hat{e}_{ij},$
where the subscript ‘.’ denotes an average over that index. The terms $\hat{e}_{ij}$ can be used to estimate the variance $\sigma^{2}_{\epsilon_{ij}}$ through the alpha model, which uses a logit transformation and may involve a rich set of covariates different from $\mathbf{X}$ in the mean regression model.
The $k$th set of synthetic values is then generated for all population units as
$y^{syn(k)}_{ij}=\hat{\beta}_{0}+\mathbf{x}_{ij}^{\prime}\hat{\bm{\beta}}+\hat{\gamma}_{i}^{(k)}+\hat{e}_{ij}^{*(k)},$
where $\hat{\gamma}_{i}^{(k)}$ is sampled with replacement from the set of $\hat{\gamma}_{i}$, and $\hat{e}_{ij}^{*(k)}$ is sampled with replacement from the set of standardized residuals and then scaled by $\hat{\sigma}_{\epsilon ij}$. The ELL estimator of the poverty indicator $F_{\alpha i}$ is finally computed as
$\hat{F}^{ELL}_{\alpha i}=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}\frac{1}{K}\sum_{k=1}^{K}h_{\alpha}(y^{syn(k)}_{ij})$.
It is important to clarify that this implementation does not aim to replicate the full applied modeling strategy described in Elbers et al. (2003), which typically involves extensive variable selection, higher-order terms, interaction effects, contextual variables, and possibly cluster-level effects. Instead, consistent with Molina and Rao (2010), we adopt a simplified structure in order to compare prediction mechanisms under a common and controlled data-generating process. This allows us to isolate differences attributable to the prediction methodology rather than to model selection or specification choices. For additional details, we refer readers to Elbers et al. (2003).
6.1 Model-based simulation I
In the first simulation study, we assume all 374 communes are selected. Since all areas are sampled in this setting (i.e., $n_{i}>0$ for all $i$), the area-specific tuning parameters $\tau_{i}$ can be estimated directly using within-area sample information, as described in Section 4.2.
Within each commune, 3% of units are selected by simple random sampling without replacement. Consequently, the within-area sample sizes, $n_{i}$, vary across communes from 4 to 2693. The welfare variable $E_{ij}$ is obtained as an exponential function of $Y_{ij}$. That is, the transformation $T(\cdot)$ is defined as $Y_{ij}=\log(E_{ij})$. The values of $Y_{ij}$ are generated according to the following model:
$Y_{ij}=9.5+\beta_{1i}x_{1ij}+\beta_{2i}x_{2ij}+\gamma_{i}+\epsilon_{ij},~i=1,\cdots,m;~j=1,\cdots,N_{i}.$
The slopes $(\beta_{1i},\beta_{2i})$, the random area effects $\{\gamma_{i}\}$ and sampling errors $\{\epsilon_{ij}\}$ are independently generated under the following three simulation conditions:
(a) Homogeneous model $(0,0):$
All areas share the same slopes $(\beta_{1i},\beta_{2i})=(-0.3,0.5)$, $\gamma_{i}\sim\mathcal{N}(0,0.1)$, and $\epsilon_{ij}\sim\mathcal{N}(0,0.5)$. This setup corresponds to the NER model considered in Molina and Rao (2010).
(b) Varying slopes $(\beta,0):$
Areas $i=1,\dots,184$ have $(\beta_{1i},\beta_{2i})=(-0.3,0.5)$, while areas $i=185,\dots,374$ have $(\beta_{1i},\beta_{2i})=(-0.5,-0.3)$. Random effects and error terms are generated as in (a).
(c) Varying slopes and heteroskedasticity $(\beta,\sigma^{2}_{\epsilon}):$
Same slope variation and random effects as in (b), but with $\epsilon_{ij}\sim\mathcal{N}(0,\sigma^{2}_{\epsilon i})$ where $\sigma^{2}_{\epsilon i}\sim\text{Gamma}(1,2)$ independently across areas.
Each scenario is independently simulated 1000 times. The performance of the above predictors is assessed using the following three criteria:
(a) Relative Bias (RB):
@@ 6.1 Model-based simulation I
The average RB across all small areas. For a given area, RB is defined as the ratio of the average difference between the estimated and true small-area FGT measures to the true small-area FGT measure, averaged over simulations.
(b) Relative Root Mean Squared Prediction Error (RRMSPE):
@@
The average RRMSPE across all small areas. For a given area, RRMSPE is the ratio of the square root of the average squared difference between the estimated and true small-area FGT measures to the true measure, averaged over simulations.
(c) Efficiency (EFF):
Defined as the ratio of the RRMSPE of each estimator or predictor to that of the baseline estimator, with MRE serving as the baseline.
Figure 3 presents the performance of our proposed algorithm in estimating the regression coefficients and variance components under the NERHDP model. Across all three simulation scenarios, no evident bias is observed in the estimation of regression coefficients. In scenarios $(0,0)$ and $(\beta,0)$ where the sampling variance is constant across areas, the proposed algorithm exhibits slight underestimation of the area-specific sampling variance.
Figure 3: Box-plots displaying ratios of estimates of regression coefficients and area specific sampling variances to their corresponding true values under repeated sampling in our model-based simulation experiment.
Table 2 presents the average values of RB, RRMSPE, and EFF for different simulation scenarios and predictors. The results show that the Monte Carlo approximations MR and CLS generally have slightly larger RB and RRMSPE compared to MRE and CLSE, respectively, but their overall performance is similar. Therefore, in the remainder of this section, we focus on evaluating the Monte Carlo approximations, as they are applicable to more general poverty indicators. According to Table 2, our proposed method, CLS, achieves smaller RB than both MR and SELL across all three simulation scenarios. As expected, in scenario $(0,0)$, i.e., the traditional nested error regression model, MR performs best in terms of RRMSPE. In contrast, for the other two scenarios, $(\beta,0)$ and $(\beta,\sigma^{2}_{\epsilon})$ where either the slopes or both the slopes and sampling variances vary across areas, CLS outperforms the other predictors in terms of RRMSPE.
Table 2: Model-based simulation I results: performance of predictors of small area FGT measures for all sampled areas
Predictor Results (%) for the following scenarios
Headcount Ratio Poverty Gap
(0, 0) ($\beta$, 0) ($\beta$, $\sigma_{\epsilon}^{2}$) (0, 0) ($\beta$, 0) ($\beta$, $\sigma_{\epsilon}^{2}$)
Average Relative Bias
DIRECT -0.023 -0.018 0.018 -0.025 -0.008 -0.007
MR 2.154 14.085 18.471 3.445 22.785 23.414
MRE 2.032 13.929 18.321 3.299 22.566 23.221
CLS 0.067 3.611 7.530 -0.648 5.939 7.297
CLSE -0.040 3.483 7.412 -0.776 5.768 7.154
SELL 15.428 94.413 93.786 25.727 179.278 148.935
Average RRMSPE
DIRECT 25.724 32.028 32.565 31.901 39.374 40.162
MR 14.543 26.096 29.419 18.875 41.540 38.479
MRE 14.453 25.955 29.287 19.069 41.226 38.197
CLS 14.967 19.908 22.095 19.069 27.943 28.318
CLSE 14.897 19.825 22.015 18.986 27.813 28.176
SELL 54.253 179.599 163.272 77.682 334.468 242.062
6.2 Model-based simulation II
In this simulation study, we intentionally exclude 161 areas from the sample, following the sampling pattern of the 2002 Albanian LSMS data, to assess the performance of our proposed EBP methods for out-of-sample areas. Specifically, 213 communes are selected, and this selection is fixed while simple random samples (without replacement) are repeatedly drawn from these areas. Consequently, 161 areas are excluded from the sampling process.
As described in subsection 4.2, we propose a model that relates $\tau_{i}$ to the auxiliary information $\bar{\bm{Z}}_{i}$. However, since $\tau_{i}$ serves as a tuning parameter, it is difficult to incorporate it directly into the population generation model. Therefore, we assume that $(\beta_{1i},\beta_{2i})$ are functions of $\bar{\bm{Z}}_{i}=(\bar{X}_{1i},\bar{X}_{2i})$, which may indirectly capture the relationship between $\tau_{i}$ and $\bar{\bm{Z}}_{i}$. In this experiment, we consider the following three data-generation scenarios:
(a) Homogeneous model $(0,0):$
All areas share the same slopes $(\beta_{1i},\beta_{2i})=(-0.3,0.5)$, $\gamma_{i}\sim\mathcal{N}(0,0.1)$, and $\epsilon_{ij}\sim\mathcal{N}(0,0.5)$.
(b) Varying slopes $(\beta,0):$
$(\beta_{1i},\beta_{2i})$ are generated through functions of $(\bar{X}_{1i},\bar{X}_{2i})$: $\beta_{1i}=0.2\times\bar{X}_{1i}$, $\beta_{2i}=5\times\bar{X}_{2i}$, where $i=1,\cdots,374$, with $\bar{X}_{1i}$ having a mean of 4.45 and standard deviation of 0.526, and $\bar{X}_{2i}$ having a mean of 0.22 and standard deviation of 0.277. The $\gamma_{i}$ and $\epsilon_{ij}$ are generated as in (a).
(c) Varying slopes and heteroskedasticity $(\beta,\sigma^{2}_{\epsilon}):$
Same slope variation and random effects as in (b), but with $\epsilon_{ij}\sim\mathcal{N}(0,\sigma^{2}_{\epsilon i})$ where $\sigma^{2}_{\epsilon i}\sim\text{Gamma}(1,2)$ independently across areas.
Tables 3 and 4 report the average values of RB, RRMSE, and EFF for various predictors when out-of-sample areas are involved. While both MR and CLS exhibit better overall performance than SELL, all three predictors show similar performance for out-of-sample areas in the scenario $(0,0)$, as expected, since both MR and CLS reduce to synthetic predictions in this case. For the scenarios $(\beta,0)$ and $(\beta,\sigma^{2}_{\epsilon})$, our proposed CLS method demonstrates the best performance in terms of RB and RRMSE, both overall and for out-of-sample areas.